OpenStudy (anonymous):

I need help with part 3.. Please help<3 Consider the line that passes through the point (-3, 2) and has a slope of 3. Part 1: Write the equation of this line using point-slope form. (2 points) Part 2: Using your equation from part 1, rewrite this equation in slope-intercept form. Make sure to show all of your work. (2 points) Part 3: Using your equation from part 2, rewrite this equation in standard form. Make sure to show all of your work. (2 points)

OpenStudy (e.cociuba):

Part 1: Plug into point slope formula: y-y1=m(x-x1) Part 2: Solve for y Part 3: Rearrange to standard form

OpenStudy (anonymous):

y - 2 = 3(x - -3) y = MX + B y = 3x + 2 Thats what I have so far.

OpenStudy (e.cociuba):

Thats correct. Great Job.

OpenStudy (anonymous):

But how do I go back to standard?

OpenStudy (e.cociuba):

The number with the x is the slope and the other number is the y-intercept.

OpenStudy (anonymous):

Im still confused.

OpenStudy (e.cociuba):

Part 3 4) use the slope and y-intercept of the above equation (part 2) 5) the standard from would look like y-a=x+b

OpenStudy (whpalmer4):

Standard form for a line is \[Ax + By = C\] where \(A, B, C\) are integers and \(A\gt0\).

OpenStudy (anonymous):

your standard form would be -3x+y = 2

OpenStudy (whpalmer4):

@kayy15 Most texts I've seen have the condition that A > 0, in which case you would want to multiply that equation by -1 to change the signs.